# Jaime Angulo Pava  and  Alexander Muñoz

### v3

@@ 2.1. Notation
$\vec{\textbf{U}}:=\begin{pmatrix}\phi(-L)\\ \phi(L)\\ \psi_{1}(L)\\ \vdots\\ \psi_{N}(L)\end{pmatrix},\ \vec{\textbf{U}}^{\prime}:=\begin{pmatrix}\phi^{\prime}(-L)\\ \phi^{\prime}(L)\\ \psi_{1}^{\prime}(L)\\ \vdots\\ \psi_{N}^{\prime}(L)\end{pmatrix}\mbox{ and }\vec{\textbf{U}}^{\prime\prime}:=\begin{pmatrix}\phi^{\prime\prime}(-L)\\ \phi^{\prime\prime}(L)\\ \psi_{1}^{\prime\prime}(L)\\ \vdots\\ \psi_{N}^{\prime\prime}(L)\end{pmatrix}.$
@@ 2.3. Operators on Krein Spaces
    $D(\mathcal{L}^{\#}):=\{y\in\mathcal{K}_{2}\mid\exists z\in\mathcal{K}_{1}:\left(\mathcal{L}x\mid y\right)_{\mathcal{K}_{2}}=\left(x\mid z\right)_{\mathcal{K}_{1}}\mbox{ for all }x\in D(\mathcal{L})\}\mbox{ and }\mathcal{L}^{\#}y=z.$ (2.5)
@@
    $\mathcal{L}$ is called $(\mathcal{K}_{1},\mathcal{K}_{2})$-unitary if $D(\mathcal{L})$ and $R(\mathcal{L})$ are dense, $\mathcal{L}$ is injective and #$\mathcal{L}^{\#}=\mathcal{L}^{-1}$.
@@ 3. The free Airy operator and deficiency indices
where $\alpha_{j}$ and $\beta_{j}$ are real numbers for any $j$. For $\mathcal{G}$ determined by the edges $e_{0}=[-L,L]$ and $e_{j}=[L,\infty)$ considering $-L$ as a terminal vertex, we understand the action of $A_{0}$ over the edges in a component-wise manner, that is, for $\textbf{U}=(\phi,\psi_{1},\dots,\psi_{N})$ posed on $\mathcal{G}$ we denote
@@
[…]
@@
From the definition of $A_{0}$ in (3.1) we have that solutions of $-i\alpha_{e}\partial_{x}^{3}u_{e}-i\beta_{e}\partial_{x}u_{e}\pm iu_{e}=0$ without any boundary condition are obtained as linear combinations of complex exponentials of the form
@@
We note that $e^{rx}$ belong to $L^{2}(-L,L)$ independent on $r$, which means $d_{-_{0}}=d_{+_{0}}=3$ along $e_{0}$. On the other hand, on a half-line $[L,\infty)$ we note $e^{rx}$ is square integrable only if the real part of $r$ is negative, which depends on the signs of $\alpha_{e}$ and $\beta_{e}$. We summarize the behaviour (on a single edge $[L,\infty)$) in the following table.
Signs of $\alpha_{e}$ and $\beta_{e}$ $d_{-_{e}}$ $d_{+_{e}}$
$\alpha_{e}>0,\beta_{e}\geqslant 0$ 2 1
$\alpha_{e}>0,\beta_{e}<0$ 2 1
$\alpha_{e}<0,\beta_{e}\geqslant 0$ 1 2
$\alpha_{e}<0,\beta_{e}<0$ 1 2
From the Von Neumann and Krein theory, one can expect self-adjoint extensions of $iA_{0}$ only if the deficiency indices agree, that is, $d_{+}=d_{-}$. In such case, according to Theorem 2.3, there exist extensions that generate unitary dynamics. Several combinations of signs of $\alpha_{e}$ and $\beta_{e}$ along the edges of $\mathcal{G}$ may lead to equal deficiency indices. For instance, in case the number of attached half-lines is $N=2k$ for some integer $k$ with the conditions $\alpha_{j},\beta_{j}>0$ if $j$ is odd and $\alpha_{j},\beta_{j}<0$ if $j$ is even, we would get deficiency indices $d_{+}$ and $d_{-}$ equal to $33k+$ 3. Thus, in this case, we know that there must be a $9(k+1)^{2}$ family of unitary groups for the case of the looping-edge graph, each one having its own representation.
@@ 4.1. Boundary system for even amount of half-lines
Let us first consider $\mathcal{G}$ consisting of an even amount $2k$ of half-lines attached to the vertex in which exactly half of them have a leading coefficient $\alpha_{e}>0$. More precisely, we consider $\mathcal{G}$ defined trough $V=\{L,-L\}$ and $E=\{[-L,L]\}\cup E_{+}\cup E_{-}$ where $|E_{-}|=|E_{+}|=k\in\mathbb{Z}^{+}$ and $e\in E_{\pm}$ if and only if ±$\pm\alpha_{e}>0$.
@@
with [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] being the $L^{2}(\mathcal{G})$-inner product in (3.5).
@@
[…]
@@
+=$\mathcal{B}_{+}=\Big{\{}\begin{pmatrix}\partial\phi(-L)\\ \partial\psi_{odd}(L)\end{pmatrix}:(\phi,\psi_{1}\dots,\psi_{N})\in H^{3}(\mathcal{G})\Big{\}}$ (4.4)
@@
=$\mathcal{B}_{-}=\Big{\{}\begin{pmatrix}\partial\phi(L)\\ \partial\psi_{even}(L)\end{pmatrix}:(\phi,\psi_{1}\dots,\psi_{N})\in H^{3}(\mathcal{G})\Big{\}}.$ (4.5)
@@
$\left(x\mid y\right)_{\mathcal{B}_{+}}:=\left(x\mid y\right)_{+}=\left(B_{+}x\mid y\right)\mbox{ and }\ \left(x\mid y\right)_{\mathcal{B}_{-}}:=\left(x\mid y\right)_{-}:=\left(B_{-}x\mid y\right),$ (4.6)
@@
$B_{+}:=\begin{pmatrix}B_{0}\\ &B_{1}\\ &&B_{3}\\ &&&\ddots\\ &&&&B_{2k-1}\end{pmatrix}\mbox{ and }B_{-}:=\begin{pmatrix}B_{0}\\ &-B_{2}\\ &&-B_{4}\\ &&&\ddots\\ &&&&-B_{2k}\end{pmatrix}.$
$(\mathcal{B}_{\pm},\left(\cdot\mid\cdot\right)_{\pm})$ are Krein spaces.
@@
$[A_{0}^{*}\textbf{U},\textbf{V}]+[\textbf{U},A_{0}^{*}\textbf{V}]=\left(\begin{pmatrix}\partial\phi(-L)\\ \partial\psi_{odd}(L)\end{pmatrix}\mid\begin{pmatrix}\partial\tilde{\phi}(-L)\\ \partial\tilde{\psi}_{odd}(L)\end{pmatrix}\right)_{+}-\left(\begin{pmatrix}\partial\phi(L)\\ \partial\psi_{even}(L)\end{pmatrix}\mid\begin{pmatrix}\partial\tilde{\phi}(L)\\ \partial\tilde{\psi}_{even}(L)\end{pmatrix}\right)_{-}.$ (4.7)
@@
$\begin{split}w:&\ (\mathcal{B}_{+}\oplus\mathcal{B}_{-})\times(\mathcal{B}_{+}\oplus\mathcal{B}_{-})\to\mathbb{C}\\ &((x,y),(\tilde{x},\tilde{y}))\mapsto\left(x\mid\tilde{x}\right)_{+}-\left(y\mid\tilde{y}\right)_{-}.\end{split}$ (4.9)
@@
By [44] we can now characterize skew-self-adjoint extensions $A$ of $A_{0}$ i.e., skew-self-adjoint restrictions of $A_{0}^{*}$ and therefore self-adjoint extensions of $iA_{0}$. Note that obviously $G(A)=\{(x,Ax):x\in D(A)\}\subset G(A_{0}^{*})$
@@ 4.1.1. Extensions generating unitary dynamics
Let $X\subset\mathcal{B}_{+}\oplus\mathcal{B}_{-}$. $X$ is $w$-self-orthogonal if and only if there is a $(\mathcal{B}_{+},\mathcal{B}_{-})$-unitary operator $\mathcal{L}$ such that $X=G(\mathcal{L})$.
@@
Let us now see #$\mathcal{L}^{\#}=L^{-1}$. First note for $x\in D(\mathcal{L})$ and $y\in R(\mathcal{L})$, from (4.11), we have
@@
The latter implies $R(\mathcal{L})\subset D(\mathcal{L}^{\#})$ with #$\mathcal{L}^{\#}y=\mathcal{L}^{-1}y$; i.e., $\mathcal{L}^{-1}\subset\mathcal{L}^{\#}$. On the other hand, if $y\in D(\mathcal{L}^{\#})$ with #$\mathcal{L}^{\#}y=x$ we have
$(\mathcal{L}z\mid y)_{-}=(z\mid x)_{+}\ \ \mbox{for all }z\in D(\mathcal{L}).$
@@ 4.1.2. Examples
Let us first consider $k=1$, i.e., two half-lines attached to a common vertex. Assume $-\alpha_{2}=\alpha_{1}=\alpha_{0}$ and $-\beta_{2}=\beta_{1}=\beta_{0}$. In particular,
@@
Since #$\mathcal{L}^{\#}=B_{+}^{-1}\mathcal{L}^{*}B_{-}=\mathcal{L}^{-1}$, it can be seen that $\mathcal{L}$ defines a $(\mathcal{B}_{+},\mathcal{B}_{-})$-unitary operator. According to Proposition 4.2 and Theorem 4.3, $\mathcal{L}$ defines a skew-self-adjoint extension $A_{\mathcal{L}}$ of $A_{0}$ defined via the relation $F(G(A_{\mathcal{L}}))=G(\mathcal{L})$. That is, an extension with domain $D(A_{\mathcal{L}})$ consisting of the functions $\textbf{V}=(\phi,\psi_{1},\psi_{2})\in D(A_{0}^{*})$ such that
@@
$-L$
$L$
$\phi$
$\psi_{1}$
$\psi_{2}$
$-L$
$L$
$\phi$
$\psi_{1}$
$\psi_{2}$
Figure 2. Examples of profiles satisfying (4.15) (on the left) and (4.16) (on the right).
More generally, consider $k\in\mathbb{Z}^{+}$ arbitrary (i.e. $2k$ half-lines attached to the common vertex $\nu=L$) and a parameter $z\in\mathbb{R}$. Under the analogous assumption $-\alpha_{2j}=\alpha_{2j-1}=\alpha_{0}$, $-\beta_{2j}=\beta_{2j-1}=\beta_{0}$ for all $j\in\{1,\dots,k\}$, so we obtain in particular for $k=2$
@@
$\mathcal{L_{\delta_{Z}}}:=\begin{pmatrix}\delta_{z}\\ &\delta_{z}\\ &&\ddots\\ &&&\delta_{z}\end{pmatrix}\mbox{ where }\delta_{z}:=\begin{pmatrix}1&0&0\\ z&1&0\\ \frac{z^{2}}{2}&z&1\end{pmatrix}.$ (4.18)
@@
$D_{\delta_{z}}:=\left\{\textbf{V}\in H^{3}(\mathcal{G})\ \mid\delta_{z}\partial\phi(-L)=\partial\phi(L)\mbox{ and }\delta_{z}\partial\psi_{2j-1}(L)=\partial\psi_{2j}(L),\ j=1,\dots,k.\right\}$ (4.19)
Note from (4.19) we obtained ”pair-wise” the conditions in (4.15). More precisely, abusing notation to identify $\psi_{-1}$ with $\phi$ at $-L$ and $\psi_{0}$ with $\phi$ at $L$ we have for any $j=0,\dots,k$ that
@@
For the following examples we assume $|E_{+}|=|E_{-}|=1$. From the deficiency indices in Section 3 and the Von Neumann and Krein theory we know the skew-self-adjoint extensions of $A_{0}$ are parametrized by $366^{2}=$ 36 parameters. We exhibit a $2$ and a $4$-parameter independent families of unitary operators that maintain the looping-edge structure (i.e. the loop condition $\phi(-L)=\phi(L)$) and are not necessarily interacting pair-wise.
@@
$\begin{split}&\phi(-L)=\phi(L),\ \psi_{1}(L)=\psi_{2}(L),\ \phi^{\prime}(-L)=\psi^{\prime}_{2}(L),\ \phi^{\prime}(-L)-\psi^{\prime}_{1}(L)=m_{1}\psi_{1}(L),\\ &\phi^{\prime\prime}(L)-\phi^{\prime\prime}(-L)=m_{2}\psi_{1}(L)\ \mbox{and }\psi_{2}^{\prime\prime}(L)-\psi_{2}^{\prime\prime}(-L)=\frac{m_{1}^{2}}{2}\psi_{1}(L)+m_{1}\psi_{1}^{\prime}(L)-m_{2}\phi(-L).\end{split}$
@@
Consider again $k=1$, $-\alpha_{2}=\alpha_{1}=\alpha_{0}$ and $-\beta_{2}=\beta_{1}=\beta_{0}$. It can be seen for $m_{1},\dots,m_{4}\in\mathbb{R}$ that
@@
$\begin{split}&\phi(L)=\phi(-L),\ \psi_{1}(L)=\psi_{2}(L),\ \phi^{\prime}(L)-\phi^{\prime}(-L)=m_{1}\phi(-L)+m_{2}\psi_{1}(L),\\ &\psi_{2}^{\prime}(L)-\phi^{\prime}(-L)=m_{4}\psi_{1}(L)+\phi(-L),\\ &\phi^{\prime\prime}(L)-\phi^{\prime\prime}(-L)=\frac{m_{1}^{2}+1}{2}\phi(-L)+m_{3}\psi_{1}(L)+\phi^{\prime}(-L)+m_{1}\psi_{1}^{\prime}(L),\mbox{and}\\ &\psi_{2}^{\prime\prime}(L)-\psi_{1}^{\prime\prime}(L)=(m_{4}-m_{3}+m_{1}m_{2})\phi(-L)+\frac{m_{2}^{2}+m_{4}^{2}}{2}\psi_{1}(L)+m_{4}\phi^{\prime}(L)+m_{2}\psi_{1}^{\prime}(L).\end{split}$

### v4

@@ 2.1. Notation
$\begin{split}\ &\vec{\mathbf{U}}:=\vec{\mathbf{U}}(L)=(\phi(-L),\ \phi(L),\ \psi_{1}(L),\ \dots,\psi_{N}(L))^{T},\\ \ &\vec{\mathbf{U}}^{\prime}:=\vec{\mathbf{U}}^{\prime}(L):=(\phi^{\prime}(-L),\ \phi^{\prime}(L),\ \psi_{1}^{\prime}(L),\ \dots,\psi_{N}^{\prime}(L))^{T},\\ \ &\vec{\mathbf{U}}^{\prime\prime}:=\vec{\mathbf{U}}^{\prime\prime}(L):=(\phi^{\prime\prime}(-L),\ \phi^{\prime\prime}(L),\ \psi_{1}^{\prime\prime}(L),\ \dots,\psi_{N}^{\prime\prime}(L))^{T}.\end{split}$
@@ 2.3. Operators on Krein Spaces
[…]
@@
    $\mathcal{L}$ is called $(\mathcal{K}_{1},\mathcal{K}_{2})$-unitary if $D(\mathcal{L})$ and $R(\mathcal{L})$ are dense, $\mathcal{L}$ is injective and $\mathcal{L}^{\#}=\mathcal{L}^{-1}$.
@@ 3. The free Airy operator and deficiency indices
where $\alpha_{j}$ and $\beta_{j}$ are real numbers for any $j$. For $\mathcal{G}$ determined by the edges $e_{0}=[-L,L]$ and $e_{j}=[L,\infty)$ considering L-L as a terminal vertex, we understand the action of $A_{0}$ over the edges in a component-wise manner, that is, for $\textbf{U}=(\phi,\psi_{1},\dots,\psi_{N})$ posed on $\mathcal{G}$ we denote
@@
$\textbf{U},\textbf{V}]:=[(\phi,\psi_{1},\dots,\psi_{N}),(\tilde{\phi},\tilde{\psi}_{1},\dots,\tilde{\psi}_{N})]_{L^{2}(\mathcal{G})}=\int_{-L}^{L}\phi\overline{\tilde{\phi}}dx+\sum_{j=1}^{N}\int_{L}^{\infty}\psi_{j}\overline{\tilde{\psi}_{j}}dx.$ (3.5)
@@
From the definition of $A_{0}$ in (3.1) we have that solutions of $i\alpha_{e}\partial_{x}^{3}u_{e}-i\beta_{e}\partial_{x}u_{e}\pm iu_{e}=0$ without any boundary condition are obtained as linear combinations of complex exponentials of the form
@@
We note that $e^{rx}$ belong to $L^{2}(-L,L)$ independent on $r$, which means $d_{-_{0}}=d_{+_{0}}=3$ along $e_{0}$. On the other hand, on a half-line $L,\infty)$ we note $e^{rx}$ is square integrable only if the real part of $r$ is negative, which depends on the signs of $\alpha_{e}$. We summarize the behavior (on a single edge $L,\infty)$) in the following table.
Sign of $\alpha_{e}$ $d_{-_{e}}$ $d_{+_{e}}$
$\alpha_{e}>0$ 2 1
$\alpha_{e}<0$ 1 2
From the Von Neumann and Krein theory, one can expect self-adjoint extensions of $iA_{0}$ only if the deficiency indices agree, that is, $d_{+}=d_{-}$. In such case, according to Theorem 2.3, there exist extensions that generate unitary dynamics. Several combinations of signs of $\alpha_{e}$ and $\beta_{e}$ along the edges of $\mathcal{G}$ may lead to equal deficiency indices. For instance, in case the number of attached half-lines is $N=2k$ for some integer $k$ with the conditions $\alpha_{j},\beta_{j}>0$ if $j$ is odd and $\alpha_{j},\beta_{j}<0$ if $j$ is even, we would get deficiency indices $d_{+}$ and $d_{-}$ equal to $3k+3$. Thus, in this case, we know that there must be a $9(k+1)^{2}$ family of unitary groups for the case of the looping-edge graph, each one having its own representation.
@@ 4.1. Boundary system for even amount of half-lines
Let us first consider $\mathcal{G}$ consisting of an even amount $2k$ of half-lines attached to the vertex in which exactly half of them have a leading coefficient $\alpha_{e}>0$. More precisely, we consider $\mathcal{G}$ identified trough $V=\{L,-L\}$ and $E=\{[-L,L]\}\cup E_{+}\cup E_{-}$ where $|E_{-}|=|E_{+}|=k\in\mathbb{Z}^{+}$ and $e\in E_{\pm}$ if and only if $\pm\alpha_{e}>0$.
@@
with $\cdot$$\cdot$] being the $L^{2}(\mathcal{G})$-inner product in (3.5).
@@
and $(\cdot\mid\cdot)$ represents the standard inner product in $\mathbb{C}^{3}$.
@@
$\mathcal{B}_{+}=\Big\{\begin{pmatrix}\partial\phi(-L)\\ \partial\psi_{odd}(L)\end{pmatrix}:(\phi,\psi_{1}\dots,\psi_{N})\in H^{3}(\mathcal{G})\Big\}$ (4.4)
@@
$\mathcal{B}_{-}=\Big\{\begin{pmatrix}\partial\phi(L)\\ \partial\psi_{even}(L)\end{pmatrix}:(\phi,\psi_{1}\dots,\psi_{N})\in H^{3}(\mathcal{G})\Big\}.$ (4.5)
@@
where $(\cdot\mid\cdot)$ represents the standard inner product in $\mathbb{C}^{3(k+1)}$ and
$B_{+}:=\operatorname{diag}(B_{0},B_{1},B_{3},\dots,B_{2k-1})\ \mbox{ and }\ B_{-}:=\operatorname{diag}(B_{0},-B_{2},-B_{4},\dots,-B_{2k}).$
(±,()±)($\mathcal{B}_{\pm}$$\left$($\cdot\mid\cdot\right$$_{\pm})$ are Krein spaces.
@@
$A_{0}^{*}\textbf{U},\textbf{V}]+[\textbf{U},A_{0}^{*}\textbf{V}]=\left(\begin{pmatrix}\partial\phi(-L)\\ \partial\psi_{odd}(L)\end{pmatrix}\mid\begin{pmatrix}\partial\tilde{\phi}(-L)\\ \partial\tilde{\psi}_{odd}(L)\end{pmatrix}\right)_{+}-\left(\begin{pmatrix}\partial\phi(L)\\ \partial\psi_{even}(L)\end{pmatrix}\mid\begin{pmatrix}\partial\tilde{\phi}(L)\\ \partial\tilde{\psi}_{even}(L)\end{pmatrix}\right)_{-}$. (4.7)
@@
[…]
@@
By [49] we can now characterize skew-self-adjoint extensions $A$ of $A_{0}$ - i.e., skew-self-adjoint restrictions of $A_{0}^{*}$ Ð and therefore self-adjoint extensions of $iA_{0}$. Note that obviously $G(A)=\{(x,Ax):x\in D(A)\}\subset G(A_{0}^{*})$
@@ 4.1.1. Extensions generating unitary dynamics
Let $\subset\mathcal{B}_{+}\oplus\mathcal{B}_{-}. XX$ is $w$-self-orthogonal if and only if there is a $(\mathcal{B}_{+},\mathcal{B}_{-})$-unitary operator $\mathcal{L}$ such that $X=G(\mathcal{L})$.
@@
Let us now see $\mathcal{L}^{\#}=L^{-1}$. First note for $x\in D(\mathcal{L})$ and $y\in R(\mathcal{L})$, from (4.11), we have
@@
The latter implies $R(\mathcal{L})\subset D(\mathcal{L}^{\#})$ with $\mathcal{L}^{\#}y=\mathcal{L}^{-1}y$; i.e., $\mathcal{L}^{-1}\subset\mathcal{L}^{\#}$. On the other hand, if $y\in D(\mathcal{L}^{\#})$ with $\mathcal{L}^{\#}y=x$ we have
[…]
@@ 4.1.2. Examples
Let us first consider $k=1$, i.e., two half-lines attached to a common vertex. Assume $\alpha_{2}=\alpha_{1}=\alpha_{0}$ and $\beta_{2}=\beta_{1}=\beta_{0}$. In particular,
@@
Since $\mathcal{L}^{\#}=B_{+}^{-1}\mathcal{L}^{*}B_{-}=\mathcal{L}^{-1}$, it can be seen that $\mathcal{L}$ defines a $(\mathcal{B}_{+},\mathcal{B}_{-})$-unitary operator. According to Proposition 4.2 and Theorem 4.3, $\mathcal{L}$ defines a skew-self-adjoint extension $A_{\mathcal{L}}$ of $A_{0}$ defined via the relation $F(G(A_{\mathcal{L}}))=G(\mathcal{L})$. That is, an extension with domain $D(A_{\mathcal{L}})$ consisting of the functions $\textbf{V}=(\phi,\psi_{1},\psi_{2})\in D(A_{0}^{*})$ such that
@@
More generally, consider $k\in\mathbb{Z}^{+}$ arbitrary (i.e. $2k$ half-lines attached to the common vertex $\nu=L$) and a parameter $z\in\mathbb{R}$. Under the analogous assumption $\alpha_{2j}=\alpha_{2j-1}=\alpha_{0}$, $\beta_{2j}=\beta_{2j-1}=\beta_{0}$ for all $j\in\{1,\dots,k\}$, so we obtain in particular for $k=2$
@@
$\mathcal{L_{\delta_{Z}}}:=\operatorname{diag}(\delta_{z},\delta_{z},\dots,\delta_{z})\ \mbox{ where }\ \delta_{z}:=\begin{pmatrix}1&0&0\\ z&1&0\\ \frac{z^{2}}{2}&z&1\end{pmatrix}$. (4.18)
@@
$D_{\delta_{z}}:=\left\{\textbf{V}\in H^{3}(\mathcal{G})\ \mid\delta_{z}\partial\phi(-L)=\partial\phi(L)\mbox{ and }\delta_{z}\partial\psi_{2j-1}(L)=\partial\psi_{2j}(L),\ j=1,\dots,k\right\}.$ (4.19)
Note from (4.19) we obtained ”pair-wise” the conditions in (4.15). More precisely, abusing notation to identify $\psi_{-1}$ with $\phi$ at L-L and $\psi_{0}$ with $\phi$ at $L$ we have for any $j=0,\dots,k$ that
@@
For the following examples we assume $|E_{+}|=|E_{-}|=1$. From the deficiency indices in Section 3 and the Von Neumann and Krein theory we know the skew-self-adjoint extensions of $A_{0}$ are parametrized by $6^{2}=36$ parameters. We exhibit a $2$ and a $4$-parameter independent families of unitary operators that maintain the looping-edge structure (i.e. the loop condition $\phi(-L)=\phi(L)$) and are not necessarily interacting pair-wise.
@@
$\begin{split}&\phi(-L)=\phi(L),\ \psi_{1}(L)=\psi_{2}(L),\ \phi^{\prime}(-L)=\psi^{\prime}_{2}(L),\ \phi^{\prime}(L)-\psi^{\prime}_{1}(L)=m_{1}\psi_{1}(L),\\ &\phi^{\prime\prime}(L)-\phi^{\prime\prime}(-L)=m_{2}\psi_{1}(L)\ \mbox{and }\psi_{2}^{\prime\prime}(L)-\psi_{2}^{\prime\prime}(L)=\frac{m_{1}^{2}}{2}\psi_{1}(L)+m_{1}\psi_{1}^{\prime}(L)-m_{2}\phi(-L).\end{split}$
@@
Consider again $k=1$, $\alpha_{2}=\alpha_{1}=\alpha_{0}$ and $\beta_{2}=\beta_{1}=\beta_{0}$. It can be seen for $m_{1},\dots,m_{4}\in\mathbb{R}$ that
@@
$\begin{split}&\phi(L)=\phi(-L),\ \psi_{1}(L)=\psi_{2}(L),\ \phi^{\prime}(L)-\psi_{1}^{\prime}(L)=m_{1}\phi(-L)+m_{2}\psi_{1}(L),\\ &\psi_{2}^{\prime}(L)-\phi^{\prime}(-L)=m_{4}\psi_{1}(L)+\phi(-L),\\ &\phi^{\prime\prime}(L)-\phi^{\prime\prime}(-L)=\frac{m_{1}^{2}+1}{2}\phi(-L)+m_{3}\psi_{1}(L)+\phi^{\prime}(-L)+m_{1}\psi_{1}^{\prime}(L),\mbox{and}\\ &\psi_{2}^{\prime\prime}(L)-\psi_{1}^{\prime\prime}(L)=(m_{4}-m_{3}+m_{1}m_{2})\phi(-L)+\frac{m_{2}^{2}+m_{4}^{2}}{2}\psi_{1}(L)+m_{4}\phi^{\prime}(-L)+m_{2}\psi_{1}^{\prime}(L).\end{split}$